# Simultaneous elections make single-party sweeps more likely

### v1

@@ 5.1 Random Voter Turnout
A key hypothesis of our model is that the probabilites of turning out to vote are independent across individuals (though they may vary arbitrarily across them). This is very much in the spirit of Penrose [30], and later Banzhaf [4]), both of whom considered the special case where these probabilities were identically $1/2$
@@
The recent “instrumental theory of turnout” [^5] The theory posits individuals for whom the act of voting is “instrumental” to maximizing their own “payoffs”, though these payoffs may be very broadly defined, incorporating not only the joy of seeing their party win but also the joy of participating in the election (alongside the cost of participating). provides theoretical underpinnings for the independence assumption. Some of the analyses in the field are focused on individual rational choice (in an exogenously fixed environment), modeled in terms of adaptive learning (e.g., [19]), or utility maximization (e.g., [39]) or minmax regret (e.g.,[12]). Others discuss full-blown strategic interaction among the voters (and sometimes also candidates) and derive Nash equilibria of “participation games” (e.g., [21], [26], [27]). These analyses differ considerably from one another on how individual turnout probabilities are formed, but they all imply the independence of those probabilities. (See the surveys in [8] and, for a broader panorama, in [16]). Some recent models (e.g., [31]), while maintaining the instrumentality assumption, introduce the possibility of communication — broadly defined — between candidates, media and voters. This gives rise to correlated equilibria where the turnout probabilities of voters are no longer independent. (Also, there are several behavioral models which openly depart from the “instrumental theory” and directly incorporate correlated turnout among the voters, such as voting in “teams” (see [11], [37] and the references therein).
@@
The independence hypothesis is crucial to our analysis. However there is a “special kind” of correlation that can be admitted in our model (which, while mathematically trivial, may be useful conceptually and in applications). As in Harsanyi ([18]) define the type of voter $h\in H$ to consist of two components: the party $t\in T$ that $h$ will vote for; and the probability $p_{h}$ with which $h$ will turn out to vote on any polling date (across all elections). Prior to the announcement of the polling schedule, let nature pick the vector of voter-types from some apriori fixed set according to a probability distribution. [^6] Two interpretations come to mind. (I) Think of nature’s “moves” as public announcement of news which affects voter behavior, increasing or decreasing their turnout probabilities and sometimes even determining which party they will vote for. (II) Suppose there is full correlation in turnout probabilities, indeed — as an extreme case — suppose even that some set of voters go to the polls with probability $1,$ while the others go with probability 00 (i.e., abstain). Our model does admit this scenario, though now the likelihood of a sweep (by any party) will be invariant of the polling schedule. However, were nature to add idiosyncratic (independent) uncertainty — no matter how small — in voters’ turnout behavior, then simultaneous polls would induce strictly higher likelihood of a sweep compared to any other polling schedule. Indeed it would suffice to add uncertainty for at least one voter $h$ who is not a “dummy” in the sense that he affects the electoral outcome with positive probability, i.e., $\chi_{l}^{t}(Z\cup\left\{h\right\})-\chi_{l}^{t}\left(Z\right)\neq 0$ for some $t,l$ and some turnout $Z$ which occurs with positive probability. For then the probability of a sweep by party $t$ would be higher under simultaneous polls than under any other schedule. Thus nature’s move is ex ante: every voter’s type stays fixed throughout the elections. Now, if schedule $\Pi$ is coarser than $\Pi^{\prime}$, then the probability of a sweep is higher under $\Pi$ than under $\Pi^{\prime}$ contingent on every move of nature (by Theorem C), therefore obviously also in expectation across all the moves of nature. Note that by allowing the type of a voter to vary, we have made some room for undecided voters who could vote for different parties. However if nature were to move ex post, repeatedly and independently between different poll dates, the argument given above breaks down.
@@ 5.2 Electoral systems
It also applies to elections based on proportional representation, where each person votes for a single party of his choice, and the percentage of parliamentary seats accorded to any party equals (as nearly as possible) the percentage of total votes it received. Of course, for our analysis to be valid, the method for “rounding” the seats (into integers) should be “monotonic”  [^7] Monotonicity holds for most of the rounding methods used in practice, e.g., the Highest Average or the Largest Remainder Methods [1] in the obvious sense, so that condition (b) is not violated; and, furthermore, condition (a) must also hold (e.g., the party with the most seats is deemed to win). Such elections are in use — with minor modifications — in many countries and also in the European Parliament (see, again, [1] for details).
@@
There is, however, a category of elections that lies outside the ambit of our model, and it includes the Two Round System (used in the French Presidential Election), Single Transferable Vote, Alternative Vote, Supplementary Vote, Borda Count and so on (described in detail in [1]). In all these cases, a vote is tantamount — either implicitly or explicitly — to a (possibly partial) ranking of the candidates/parties involved. Such elections have received the lion’s share of attention in the theoretical literature on “social choice” (SC) ever since Arrow’s pioneering monograph [3]. (For a succinct survey, see, e.g., [14].) The concerns of the SC literature are quite different from ours, but even from the “high level” viewpoint, there are salient differences. First, our voter only specifies her/his top choice instead of ranking the candidates. More importantly, in SC, the entire voter population is assumed to always turn out to vote. In sharp contrast, it is the variable, stochastic turnout of voters on which our entire analysis turns. It might be interesting to examine, along the lines of this paper, how the probability distribution of electoral outcomes depends on the stochastic turnout of voters, in the context of the more sophisticated elections considered in SC (e.g., Borda Count, see [23]), but that would be a topic for another paper.
5.3 Single-Party Sweeps
A single-party sweep (also known as “one-party monopoly” or “one-party dominance”) is a widely prevalent phenomenon, both in local and national elections ([28], [29], [38]) Its causes have been discussed from different standpoints, including: the strategic timing of elections ( [20]), economic intervention ([20]), manipulation of electoral rules ([24] ), and fragmented opposition ([40]). There does not seem to be much in the literature connecting polling schedules to single-party sweeps, which is the focus of this paper. Two notable exceptions are the “coattail effect” in the context of US elections (see, e.g., [6, 7, 25]). and a recent study ([5]) of the Indian elections. In both these cases, the analysis turns (see the next two sections) on the postulate that a voter’s behavior changes when elections are held simultaneously, on account of either the popularity of the Presidential candidate or the “salience” of one of the political parties. In sharp contrast, voters’ behavior is fixed in our model, regardless of the polling schedule or the characteristics of the candidates standing for election. It is the statistical aggregate of voters’ behavior that nevertheless varies with the schedule and drives our result. Thus our analysis brings to light a cause for single-party sweeps that is not only novel, but invariant of the candidates’ influence on voters and — to that extent — more constant.
5.4 US Elections and The Coattail Effect
The data from USA shows that a single party is more likely to win multiple elections when they are held simultaneously. This phenomenon is referred to as the “coattail” effect, based on the the rationale that a popular presidential candidate fosters additional wins for his party in Congress, with many members of Congress voted into office “on the coattails” of the president. Indeed, during Congressional elections in USA that coincide with the presidential race, the party of a popular president tends to win more seats in Congress compared to elections held in the midterm of the Presidency (the latest instance of this being the US elections in November 2024).
Our analysis brings to light a completely different mechanism for a single-party sweep, which is not at odds with the coattail effect, but independent of it. Two behavioral causes are commonly postulated for the coattail effect: a popular presidential candidate can mobilize each member in his party base to turn out to vote with higher probability; and, moreover, he can sway undecided voters to his party’s fold. [^8] A cause of a different nature is the “anti-incumbency factor” so often alluded to by political analysts. It can come into play during midterm elections, especially if there is widespread perception that the ruling party has been unable to fulfil its electoral promises. For then the rival party makes gains in the midterm elections, further accentuating the benefit of simultaneous elections for the ruling party. These causes are ruled out in our model since the party that $h$ will vote for, as well the probability $p_{h}$ that he will vote on any poll date, are both exogenously fixed for every $h\in H$ and not susceptible to change. It is purely “the statistics of voting” that causes the single-party sweep. Were the coattail effect to be present, it would further boost the probability of a sweep for the favored party [^9] We could incorporate the “coattail” in our model as an increase in the probabilities of turnout of the adherents of the favored party in the Presidential election and in (at least) all other contemporaneous Congressional elections (while the turnout probabilities of adherents of rival parties stay fixed or go down). This will obviously increase the probability of win for the favored party in every election. In this sense, our model is not opposed to the coattail effect, but accomodative of it.; but even in its absence, the likelihood of a sweep increases as the elections are made “more simultaneous” in the precise sense of Theorems B and C.
The probability of a single-party sweep can, of course, be very small if there are many parties or many elections. It becomes significant when, as is often the case, there are two parties contending two (or three) elections. To the extent that “Duverger’s Law” operates and a representative democracy settles into a two-party system, the sweep can be quite a robust phenomenon [10, 13, 32].
5.5 Indian Elections: “ONOE” & “Double-Engine” Governance
Elections in India are held as follows. Political parties contend for seats in the Lok Sabha at the national level and also in $28$ Vidhan Sabhas, one for each state. The country (resp., state) is partitioned into electoral districts for the Lok (resp., Vidhan) Sabha election, the candidate with the most votes is elected from each district, and the party with the most candidates is deemed to have won the election. [^10] W.l.o.g. we may assume that the entire electorate of India participates in each Vidhan Sabha election, viewing voters outside the state as “strategic dummies” whose votes do not count. (Condition (b) still holds for such dummies, and therefore our model applies.)
The possibility of simultaneous elections is no longer theoretical but it looms large in India and is the subject of a fierce ongoing debate, set into motion by the BJP and its allies (currently in power in the Lok Sabha) seeking a Constitutional Amendment to implement their “One Nation, One Election” (ONOE) proposal (which would make all Vidhan Sabha elections simultaneous with the Lok Sabha election). At the same time, the BJP has been strongly campaigning for “double-engine” governance, especially during elections in the states. [^11] This expression signifies the scenario in which the same party governs in both the Center and the State (for then the two can presumably work more in concert — as a “double engine” pulling in the same direction — for the development of the State). Our Theorem C shows that the likelihood of one party winning the Lok Sabha and the Vidhan Sabha of any given state is higher when both elections are held simultaneously compared to when they are not. It immediately follows that, under ONOE, the probability of “double-engine” governance would go up for every state. This would be a systemic change in Indian elections, brought on by ONOE, for it would pertain no matter which party wins the Lok Sabha.
Finally recall that the notion of “win” is very general in our model and can vary across elections, subject only to conditions (a) and (b). This enables us to consider Indian elections from diverse viewpoints, including the following two starkly different ones: $1+28=29$ “literal” elections, or $2$ “figurative” elections (with the $28$ Vidhan Sabha elections viewed as one conglomorate). The literal elections clearly fit into our model, but here the likelihood of a sweep is so miniscule that it is unlikely for either party to be focused on it. Sweeps occur with higher ptobability in the $2$-election framework, where “win” in Vidhan Sabhas can be defined in diverse ways, e.g., (i) winning a specific set of states [^12] The contending parties may have different sets of states in their sights but, so long as the intersection of their sets is non-empty, condition (a) holds and our analysis remains intact.; (ii) winning most states overall; (iii) winning most seats overall (relevant for the future Rajya Sabha); and so on. All these scenarios are covered by Theorem C, which allows for staggered polling schedules (i.e., personal partitions of the elections in $L)$.

### v2

@@ 3.1 Random Voter Turnout
A key hypothesis of our model is that the probabilites of turning out to vote are independent across individuals, and also across different polling dates for any given individual [^7] Furthermore, while these probabilities may vary arbitrarily across individuals, they remain constant (across different polling dates) for each individual.. This is very much in the spirit of Penrose [33], and later Banzhaf [5]), both of whom considered the special case where these probabilities were identically $1/2$
@@
The recent “instrumental theory of turnout” [^8] The theory posits individuals for whom the act of voting is instrumental in maximizing their own “payoffs”, though these payoffs may be very broadly defined, incorporating not only the joy of seeing their party win but also the joy of participating in the election (alongside the cost of participating). provides theoretical underpinnings for the independence assumption. Some of the analyses in the field are focused on individual rational choice (in an exogenously specified environment), modeled in terms of adaptive learning (e.g., [22]), or utility maximization (e.g., [43]) or minmax regret (e.g.,[14]). Others discuss full-blown strategic interaction among the voters (and sometimes also candidates) and derive Nash equilibria of “participation games” (e.g., [24], [29], [30]). These analyses differ considerably from one another on how individual turnout probabilities are formed, but they all imply the independence of those probabilities (see the surveys in [10] and [18]). Some recent models (e.g., [34]), while maintaining the instrumentality assumption, introduce the possibility of communication — broadly defined — between candidates, media and voters. This gives rise to correlated equilibria where the turnout probabilities of voters are no longer independent. (Also, there are several behavioral models which openly depart from the “instrumental theory” and directly incorporate correlated turnout among the voters, such as voting in “teams” (see [13], [41] and the references therein).
@@
The independence hypothesis is central to our analysis [^9] We take it to be a behavioral hypothesis which governs voters’ turnout. In other words, no matter how the behavior might once have originated from rational (optimal) choice, it has got entrenched as a “rule of thumb” for a voter (see [4]), which he follows without bothering to recompute the exact optimal choice, each time the election scenario changes on account of alterations in the polling schedule.. However there is a “special kind” of correlation that can be admitted in our model (which, while mathematically obvious, may be useful both conceptually and for applications). As in Harsanyi ([20]), define the type of voter $h\in H$ to consist of two components: the party $t\in T$ that $h$ will vote for; and the probability $p_{h}$ with which $h$ will turn out to vote on any polling date. Before the announcement of the polling schedule, assume that nature picks the vector of voter-types according to some a priori probability distribution on a given set. [^10] A “move” of nature could represent public news that affects voters’ types; or — alternatively — independendent (idiosyncratic) perturbations in every voter’s $p_{h}$. Thus nature’s move is ex ante, and every voter’s type stays fixed throughout the elections. Now, if schedule $\Pi$ is coarser than $\Pi^{\prime}$, then the probability of a sweep is higher under $\Pi$ than under $\Pi^{\prime}$ contingent on every move of nature (by Theorem C), therefore obviously also in expectation across all the moves of nature. (However, if nature were to move ex post independently between different poll dates, the argument given above breaks down.)
@@ 3.2 Electoral systems
It also applies to elections based on proportional representation, where each person votes for a single party of his choice, and the percentage of parliamentary seats accorded to any party equals (as nearly as possible) the percentage of total votes it received. Of course, for our analysis to be valid, the method for “rounding” the seats (into integers) should be “monotonic”  [^11] Monotonicity holds for most of the rounding methods used in practice, e.g., the Highest Average or the Largest Remainder Methods [1] in the obvious sense, so that condition (b) is not violated; and, furthermore, condition (a) must also hold (e.g., the party with the most seats is deemed to win; or, alternatively, if no party gets a strict majority of the seats, it is a non-win for every party). Such elections are in use — with minor modifications — in many countries and also in the European Parliament (see, again, [1] for details).
@@
There is, however, a category of elections that lies outside the ambit of our model, and it includes the Two Round System (used in the French Presidential Election), Single Transferable Vote, Alternative Vote, Supplementary Vote, Borda Count and so on (described in detail in [1]). In all these cases, a vote is tantamount to a ranking of the electoral candidates. Such elections have received the lion’s share of attention in the theoretical literature on “social choice” (SC) ever since Arrow’s pioneering monograph [3]. (For a succinct survey, see, e.g., [16].) The concerns of the SC literature are quite different from ours, but even from a “high level”, there are salient differences. First, our voter only specifies his top choice instead of ranking the candidates. More importantly, in SC, the entire voter population is assumed to always be present. In sharp contrast, it is the variable, stochastic turnout of voters on which our analysis turns. It might be interesting to examine, along the lines of this paper, how the probability distribution of electoral outcomes depends on the stochastic turnout, in the context of the more sophisticated elections considered in SC (e.g., Borda Count, see [26]), but that would be a topic for another paper.
@@ 3.3 Party Alliances
3.4 Single-Party Sweeps
A single-party sweep (also known as “one-party monopoly” or “one-party dominance”) is a widely prevalent phenomenon, both in local and national elections ([31], [32], [42]) Its causes have been discussed from different standpoints, including: the strategic timing of elections ( [23]), economic intervention ([23]), manipulation of electoral rules ([27] ), and fragmented opposition ([44]). There does not seem to be much in the literature connecting polling schedules to single-party sweeps, which is the focus of this paper. Two notable exceptions are the “coattail effect” in the context of US elections (see, e.g., [8, 9, 28]). and a recent study ([6]) of the Indian elections. Both analyses are based on the postulate that a voter’s behavior changes when elections are held simultaneously, on account of either the popularity of the Presidential candidate or the “salience” of one of the political parties (see the next two sections). In sharp contrast, voters’ behavior is fixed in our model, regardless of the polling schedule or the characteristics of the candidates standing for election. Nevertheless the voter turnout varies with the schedule and drives our result.
3.5 US Elections and The Coattail Effect
The data from USA shows that a single party is more likely to win multiple elections when they are held simultaneously. Indeed, during Congressional elections in USA that coincide with the presidential race, the party of a popular president tends to win more seats in Congress compared to elections held in the midterm of the Presidency (the latest instance of this being the US elections in November 2024). This phenomenon is referred to as the “coattail” effect. The rationale is that a popular presidential candidate fosters additional wins for his party in Congress, with many members of Congress voted into office “on the coattails” of the president.
Our analysis brings to light a completely different mechanism for a single-party sweep, which is not at odds with the coattail effect, but independent of it. Two behavioral causes are commonly postulated for the coattail effect: a popular presidential candidate can mobilize each member in his party base to turn out to vote with higher probability; and, moreover, he can sway undecided voters to his party’s fold. [^12] A cause of a different nature is the “anti-incumbency factor” so often alluded to by political analysts. It can come into play during midterm elections, especially if there is widespread perception that the ruling party has been unable to fulfil its electoral promises. For then the rival party makes gains in the midterm elections, further accentuating the benefit of simultaneous elections for the ruling party. These causes are ruled out in our model since the party that $h$ will vote for, as well the probability $p_{h}$ that he will vote on any poll date, are both exogenously fixed for every $h\in H$ and not susceptible to change. It is purely “the statistics of voting” that enhances the single-party sweep for simultaneous elections. [^13] We could incorporate the “coattail” in our model as an increase in the probabilities of turnout of the adherents of the President’s party in (at least) all other contemporaneous Congressional elections (while the turnout for its rivals stay fixed or go down). This will obviously increase the probability of win for the party in every election. In this sense, our model is accomodative of the coattail effect.
The probability of a single-party sweep can, of course, be very small if there are many parties or many elections. It becomes significant when, as is often the case, there are two parties contending two (or three) elections. To the extent that “Duverger’s Law” operates and a representative democracy settles into a two-party system, the sweep can be quite a robust phenomenon [12, 15, 35]
3.6 Indian Elections: “One Nation, One Election” and “Double Engine Governance”
Elections in India are held as follows. Political parties contend for seats in the Lok Sabha at the national level and also in $28$ Vidhan Sabhas, one for each state. The country (resp., state) is partitioned into electoral districts for the Lok (resp., Vidhan) Sabha election, the candidate with the most votes is elected from each district, and the party with the most candidates is deemed to have won the election. [^14] W.l.o.g. we may assume that the entire electorate of India participates in each Vidhan Sabha election, viewing voters outside the state as “strategic dummies” whose votes do not count (since condition (b) still holds for such dummies.)
The possibility of simultaneous elections is no longer theoretical but it looms large in India and is the subject of a fierce ongoing debate, set into motion by the BJP and its allies (currently in power in the Lok Sabha) seeking a Constitutional Amendment to implement their “One Nation, One Election” (ONOE) proposal (which would make all Vidhan Sabha elections simultaneous with the Lok Sabha election). At the same time, the BJP has been strongly campaigning for “double-engine” governance, especially during elections in the states. Theorem C shows that the likelihood of one party winning the Lok Sabha and the Vidhan Sabha of any given state is higher when both elections are simultaneous compared to when they are not, i.e., under ONOE, the probability of “double-engine” governance goes up for every state. This happens by way of a systemic change, no matter which party wins the Lok Sabha. The magnitude of the change can, in general, be quite significant [^15] We thank Vijay Vazirani for asking this question.. Indeed consider two parties, BJP and its allies versus the “INDIA” coalition of all their rivals (as in the 2024 elections) contending the Lok Sabha at the Center and the Vidhan Sabha in Bihar. For ease of computation assume that each contender is favored by half the population of India, and of Bihar; and that the turnout probability is $1/2$ for every voter. Then, by the law of large numbers, the probability of a sweep ( i.e., of double-engine governance) rises from about $50\%$ to about $100\%$ when the polling schedule of the two elections is made simultaneous (and rises from $25\%$ to $50\%$ for each contender).
Finally recall that the notion of “win” is very general in our model and can vary across elections, subject only to conditions (a) and (b). As mentioned in the introduction, this enables us to consider Indian elections from diverse viewpoints, including the following two starkly different ones: $1+28=29$ “literal” elections, or $2$ “figurative” elections (with the $28$ Vidhan Sabha elections viewed as one conglomorate). The literal elections clearly fit into our model, but here the likelihood of a sweep is so miniscule that it is unlikely for either party to be focused on it. Sweeps occur with higher ptobability in the $2$-election framework, where “win” in Vidhan Sabhas can be defined in diverse ways, e.g., (i) winning a specific set of states [^16] In particular, the set could consist of a single state, in which case a sweep reduces to the “double-engine governance” discussed earlier. More generally, in theory if not in practice, the contending parties may have different sets of states in their sights but, so long as the intersection of their sets is non-empty, condition (a) holds and our analysis remains intact. ; (ii) winning most states overall; (iii) winning most seats overall (relevant for the future Rajya Sabha); and so on. All these scenarios are covered by Theorem C, which allows for staggered polling schedules (i.e., personal partitions of the elections in $L)$.